OpenStudy (anonymous):

Integral of rad(x^4+4x^2)dx 0

OpenStudy (isaiah.feynman):

What's rad?

OpenStudy (anonymous):

radical or square root

OpenStudy (psymon):

\[\int\limits_{}^{}\sqrt{x^{4}+4x ^{2}}dx\] \[\int\limits_{}^{}\sqrt{x ^{2}(x ^{2}+4)}dx=\int\limits_{}^{}x \sqrt{x ^{2}+4}dx\] Now u -sub :3

OpenStudy (anonymous):

alrite thx!

OpenStudy (psymon):

Yep yep :3